OpenStudy (anonymous):

Eliminate the parameter. x = 5t, y = t + 8

OpenStudy (anonymous):

y = t + 8 implies t = y - 8. Substitution to x = 5t gives x = 5(y - 8).

OpenStudy (anonymous):

Well It has to be in the form of y=??? can you help with that?

OpenStudy (anonymous):

Sure. x = 5(y - 8) x = 5y - 40 -5y = -x - 40 \(y = \frac{1}{5}x+8\)

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

Your welcome! :D